A 20% dye solution is to be mixed with a 90% dye solution to get 140 liters of a 50% solution. How many liters of the 20% and 90% solutions will be needed?
Answer by macston(5194)

.
x=amount of 20% solution; y=amount of 90% solution=140L-x
.
0.20x+0.90y=0.50(140L)
0.20x+0.90(140L-x)=0.50(140L)
0.20x+126L-0.90x=70L
-0.70x=-56L
x=80L
ANSWER 1: 80 liters of 20% solution will be needed.
.
y=140L-x=140L-80L=60L
ANSWER 2: 60 liters of 90% solution will be needed.
.
CHECK:
0.20x+0.90y=0.50(140L)
0.20(80L)+0.90(60L)=0.50(140L)
16L+54L=70L
70L=70L